h(x)=2-xsquared find h(4)
Found 2 solutions by richard1234, ewatrrr:
Answer by richard1234(7193)
Answer by ewatrrr(24785)


Hi,
Finding h(4) is done by simply substituting 4 for x
h(x)= 2- x^2
h(4)= 2 - 4^2
h(4)= 2 - 16
h(4)= -14